OpenStudy (anonymous):

HELP PLEASE Consider the line -4x+9y=9 Find the equations of the line that is perpendicular & parallel to this line and passes through the points (3,-3)

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

we need to find the slope of the line given so we will put it in y = mx + b form and m is the slope. -4x + 9y = 9 -- add 4x to both sides 9y = 4x + 9 -- divide both sides by 9 y = 4/9x + 1 so the slope of this line is 4/9. you understanding this so far ?

OpenStudy (anonymous):

yes.. i got that far. i got y=4/9x+1 but then i couldnt figure out how to make it parallel and perpendicular. i tried using (3,-3) as the x & y.. but it wasnt working

OpenStudy (texaschic101):

a parallel line will have the same slope.... 4/9. use y - y1 = m(x - x1) slope(m) = 4/9 (3,-3) x1 = 3 and y1 = -3 now lets sub y - (-3) = 4/9(x - 3) y + 3 = 4/9x - 12/9 y = 4/9x - 12/9 - 3 y = 4/9x - 12/9 - 27/9 y = 4/9x - 39/9 y = 4/9x - 13/3 ==> parallel line understand so far ?

OpenStudy (anonymous):

yes!

OpenStudy (texaschic101):

a perpendicular line will have the negative reciprocal of the slope. Negative reciprocal of 4/9 is -9/4 (flip the slope and change the sign. y - y1 = m(x - x1) slope(m) = -9/4 (3,-3) x1 = 3 and y1 = -3 subbing tme y - (-3) = -9/4(x - 3) y + 3 = -9/4x + 27/4 y = -9/4x + 27/4 - 3 y = -9/4x + 27/4 - 12/4 y = -9/4x + 15/4 ==> your perpendicular line

OpenStudy (texaschic101):

any questions ?